### v1

@@ Abstract
We conduct two searches for continuous, nearly monochromatic gravitational waves originating from the central compact objects in the supernova remnants Cassiopeia A and Vela Jr. using public LIGO data. The search for Cassiopeia A targets signal frequencies between 20 Hz and 400 Hz; the Vela Jr. search between 400 Hz and 1700 Hz, and both investigate the broadest set of waveforms ever considered with highly sensitive deterministic search methods. Above 1500 Hz the Vela Jr. search is the most sensitive carried out thus far, improving on previous results by over 300%. Above 976 Hz these results improve on existing ones by 50%. In all we investigate over $10^{18}$ waveforms, leveraging the computational power donated by thousands of Einstein@Home volunteers. We perform a 4-stage follow-up on more than 6 million waveforms. None of the considered waveforms survives the follow-up scrutiny, indicating no significate detection candidate. Our null results constrain the maximum amplitude of continuous signals as a function of signal frequency from the targets. The most stringent 90% confidence upper limit for Cas A is $h_{0}^{90\%}\approx 7.3\times 10^{-26}$ near 200 Hz, and for Vela Jr. it is $h_{0}^{90\%}\approx 8.9\times 10^{-26}$ near 400 Hz. Translated into upper limits on the ellipticity and r-mode amplitude, our results probe physically interesting regions: for example the ellipticity of Vela Jr. is constrained to be smaller than $10^{-7}$ across the frequency band, with a tighter constraint of less than $2\times 10^{-8}$ at the highest frequencies.
@@ 5 The Einstein@Home Search
We employ a “stack-slide” approach for semi-coherent searches, leveraging the Global Correlation Transform (GCT) method (H. J. Pletsch, 2008; H. J. Pletsch & B. Allen, 2009; H. J. Pletsch, 2010). In this type of search, the dataset, which spans a total observing time $T_{\text{obs}}$, is partitioned into shorter $N_{\text{seg}}$ segments of equal span $T_{\text{coh}}$. The data in each segment $i$ is analyzed coherently. Afterwards, the coherent detection statistics ${\mathcal{F}}_{i}$ from all the segments are summed. The coherent detection statistic from each segment is a matched filter between the data and a waveform model with parameters {$f$, $\dot{f}$, $\ddot{f}$, $\alpha$, $\delta$}. The final detection statistic is
@@
The search setup consists of the coherent baseline time $T_{\mathrm{coh}}$, the signal-parameter ranges and the grid-spacings of the templates $\delta f$, $\delta\dot{f}$ and $\delta\ddot{f}$. For each of the searches, the grid-spacings in frequency and spindown parameters over the parameter space are constant. Every search setup has an associated average mismatch $\bar{m}$, which measures the average loss in signal-to-noise ratio due to the mismatch between the parameters of a signal and the closest ($f$,$\dot{f}$, and $\ddot{f}$) grid point. The search setups are given at the top of Table 1.
We use different search ranges for Cas A (“C”) and Vela Jr (“V”):
@@
6 Post-procession of Einstein@Home Results
@@
Similar to J. Ming et al. (2024), $R^{a}_{\mathrm{thr}}$ is determined by requiring that it is safe against discarding signals from our target population (see next Section). So thousands of fake signals from the target population are added to the data, each is searched for, exactly as done for the real search, and the corresponding $R^{a}$ is measured. We ensure that the threshold dismisses less than $0.01\%$ of the fake signals.
@@ 6.1 Target population
The post-processing procedures that follow the Einstein@Home search have been designed to have a very low false dismissal for signals of the target population. The target population signals for a given supernova remnant obviously all come from the position of that source and have frequency and frequency derivative values in the target range, randomly displaced from the search grid points. The $\cos\iota$ and $\psi$ parameters are uniformly distributed $-1\leq\cos\iota\leq 1$, $-\pi/4\leq\psi\leq\pi/4$. The signal strength is set to the target strength with respect to the amplitude spectral density: the target population signals have amplitudes corresponding to $\mathcal{D}=90~{}[1/\sqrt{\textrm{Hz}}]$ for both Vela Jr. and Cas A (see Eq. 15 for the relation between the amplitude $h_{0}$ and $\mathcal{D}$).
@@ 6.2 Stage-0
The goal at this stage is to efficiently select the most compelling candidates from the $5.7\times 10^{10}$ results returned to the Einstein@Home server. Historically, there has been three main ways to achieve this.
@@
The third method is called clustering. In clustering, candidates associated with the same astrophysical or non-astrophysical causes are identified, and one of them is chosen (A. Singh et al., 2017; B. Steltner et al., 2022a). Clustering is generally efficient because it prevents the waste of resources on following up candidates originatign from the same root cause. While clustering has proven to be very effective in all-sky searches, in directed searches the sensitivity improvement compared with simple thresholding is very modest – less than $1$ (J. Ming et al., 2024). Since determining a good operating point for the clustering procedure requires time and extensive Monte Carlos, we do not pursue it here. Instead, we propose an alternative scheme to select the Stage-0 results to follow-up: the pixeling procedure. This alternative is straightforward and efficient, and does not loose detection efficiency in disturbed frequency bands.
@@ 6.2.1 The Pixeling Procedure
We divide the searched frequency range in $N_{\textrm{bands}}$ bands of equal size. For each frequency band, we consider $N_{p}$ frequency-spindown “pixels”. We take the top $n_{px}$ Stage-0 results from waveforms with parameters in each pixel, with $n_{px}$, typically determined by the available computational budget for the follow-up.
Specifically we consider 0.5 Hz wide frequency bands, and $N_{p}=100\times 10$ pixels to cover the $f\times\dot{f}$ parameter space. We take the top $n_{px}=2$ results from each pixel. We call these candidates. The lowest detection statistic in the selected pixels from a 50 mHz region is shown in Figure 2 as an example.
In all we select $\approx 1.5$ million candidates from the Cas A search and $\approx 5.2$ million from Vela Jr. search. Both these numbers are $<2N_{p}\times N_{\textrm{bands}}$ (twice the total number of pixels) because not all pixels contribute 2 candidates – in fact some contribute no candidate at all. This is due to the fact that the original results are themselves top-lists returned by the Einstein@Home volunteer computers, and in the presence of very large disturbances most results might be concentrated in certain frequency-spindown regions rather than being $\approx$ uniformly distributed among the pixels. The average value of the $\hat{\beta}_{\text{S/GLtLr}}$ detection statistic of the considered candidates increases by 4 going from 20 to 400 Hz for Cas A and increases by 2 going from 400 to 1700 Hz for Vela Jr.. This is due to the fact that we are taking the most significant 2 results from a parameter space that increases in size, as shown in Figure 1.
@@
Figure 2: An example of the lowest recorded $\hat{\beta}_{\text{S/GLtLr}}$ per pixel in a slightly disturbed (top-left) band. The results shown come from the Vela Jr. search.
@@ 7.1 Intrinsic Gravitational Wave Amplitude Upper Limits
The detection efficiency is the number of recovered signals relative to the total number of signals considered. Each test signal is added to the data and then the with the test signals is treated exactly like the data from original Einstein@Home search, including data preparation, cleaning and candidate-selection with pixeling. A signal is considered recovered if a result from the search with the fake signal, and ascribable to the signal, survives stage-0.
As described in Section 4, in frequency bands affected by large spectral-line disturbances, we substitute the real data with Gaussian noise. For consistency we also do this after the fake signal has been added to noise, hence, depending on the data removed ands on the frequency of the fake signal, it might not be possible to reach the 90% confidence level for any value of $h_{0}$. For frequency bands where this happens we do not provide upper limits. We add these bands to the list of excluded bands given in J. Ming et al. (2025). Such bands include 6 half-Hz bands between 20 Hz and 400 Hz and 109 half-Hz bands between 400 Hz and 1700 Hz search, mostly due to the thermal excitation of vibration modes of the suspension silica fibers at 500 Hz and their harmonics around 1000 Hz and 1500 Hz (B. P. Abbott et al., 2016).
@@ 7.2 Upper Limits Recast
$\alpha=\left(\frac{h_{0}}{10^{-24}}\right)\left(\frac{D}{1\ \text{kpc}}\right)\left(\frac{f}{100\ \text{Hz}}\right)^{3}.$ (14)
@@ 7.3 Sensitivity Depth
The average sensitivity depth across the frequency ranges is:
$\begin{cases}\textrm{Cas A~{}~{} 20-400 Hz}:~{}&{{\mathcal{D}}}^{90\%}\approx 92~{}[{1/\sqrt{\text{Hz}}}]\\ \textrm{Vela Jr.~{}~{} 400-1700 Hz}:~{}&{{\mathcal{D}}}^{90\%}\approx 89~{}[{1/\sqrt{\text{Hz}}}].\end{cases}$ (16)
Although the average mismatch of the original Einstein@Home search for the Vela Jr. is slightly smaller than for Cas A, the average sensitivity depth of search for Vela Jr. is slightly larger. The reason for this is that the follow-ups involve proportionally more candidates for Vela Jr. than for Cas A: the original search for Vela Jr. has about 12 times more templates than the search of Cas A, while the number of followed-up candidates of Vela Jr. is just about 3 times more than the one from Cas A.
@@ 8 Conclusions
We present results for searches for continuous gravitational waves from the neutron star in the Cas A and Vela Jr. supernova remnants. The considered waveforms cover frequencies in the range 20 - 400 Hz for Cas A and 400 - 1700Hz Vela Jr., and very broad frequency evolution scenarios including power-las frequency evolutions with braking indexes up to 7. In particular, the second-order spin-down range of our searches is $\ddot{f}\in[0,7{f/\tau^{2}}]$, which is broader than the range covered by R. Abbott et al. (2022d) and J. Wang & K. Riles (2024), who assume $[2|\dot{f}|^{2}/f,7|\dot{f}|^{2}/f]$. Furthermore, our search range is independent of the template $\dot{f}$ value, accommodating signals that deviate from a strict power law.
The large parameter space, comprising in all $1.3\times 10^{18}$ waveforms, could be investigated thanks to the computing power donated by the volunteers of the Einstein@Home project. Promising candidates across this space were identified using a new “pixeling” method and followed-up in a cascade of further searches. Fewer and fewer candidates survive each Stage-and at the end of the fourth stage no signal candidate is left, indicating no detectable signal.

### v2

@@ Abstract
We conduct two searches for continuous, nearly monochromatic gravitational waves originating from the central compact objects in the supernova remnants Cassiopeia A and Vela Jr. using public LIGO data. The search for Cassiopeia A targets signal frequencies between 20 Hz and 400 Hz; the Vela Jr. search between 400 Hz and 1700 Hz, and both investigate the broadest set of waveforms ever considered with highly sensitive deterministic search methods. Above 1500 Hz the Vela Jr. search is the most sensitive carried out thus far, improving on previous results by over 300%. Above 976 Hz these results improve on existing ones by 50%. In all we investigate over $10^{18}$ waveforms, leveraging the computational power donated by thousands of Einstein@Home volunteers. We perform a 4-stage follow-up on more than 6 million waveforms. None of the considered waveforms survives the follow-up scrutiny, indicating no significant detection candidate. Our null results constrain the maximum amplitude of continuous signals as a function of signal frequency from the targets. The most stringent 90% confidence upper limit for Cas A is $h_{0}^{90\%}\approx 7.3\times 10^{-26}$ near 200 Hz, and for Vela Jr. it is $h_{0}^{90\%}\approx 8.9\times 10^{-26}$ near 400 Hz. Translated into upper limits on the ellipticity and r-mode amplitude, our results probe physically interesting regions: for example the ellipticity of Vela Jr. is constrained to be smaller than $10^{-7}$ across the frequency band, with a tighter constraint of less than $2\times 10^{-8}$ at the highest frequencies.
@@ 5 The Einstein@Home Search
We employ a “stack-slide” approach for semi-coherent searches (P. R. Brady et al., 1998; P. R. Brady & T. Creighton, 2000),leveraging the Global Correlation Transform (GCT) method (H. J. Pletsch, 2008; H. J. Pletsch & B. Allen, 2009; H. J. Pletsch, 2010). In this type of search, the dataset, which spans a total observing time $T_{\text{obs}}$, is partitioned into shorter $N_{\text{seg}}$ segments of equal span $T_{\text{coh}}$. The data in each segment $i$ is analyzed coherently. Afterwards, the coherent detection statistics ${\mathcal{F}}_{i}$ from all the segments are summed. The coherent detection statistic from each segment is a matched filter between the data and a waveform model with parameters {$f$, $\dot{f}$, $\ddot{f}$, $\alpha$, $\delta$}. The final detection statistic is
@@
The search setup consists of the coherent baseline time $T_{\mathrm{coh}}$, the signal-parameter ranges and the grid-spacings of the templates $\delta f$, $\delta\dot{f}$ and $\delta\ddot{f}$. For each of the searches, the grid-spacings in frequency and spindown parameters over the parameter space are constant. Every search setup has an associated average mismatch $\bar{m}$, which measures the average loss in signal-to-noise ratio due to the mismatch between the parameters of a signal and the closest ($f$, $\dot{f}$ and $\ddot{f}$) grid point. The search setups are given at the top of Table 1.
The search ranges of the spin-down parameters are frequency-dependent, as first proposed by K. Wette et al. (2008). We use different search ranges for Cas A and Vela Jr.:
@@
6 Post-processing of Einstein@Home Results
6.1 Overview
@@
Similar to J. Ming et al. (2024), $R^{a}_{\mathrm{thr}}$ is determined by requiring that it is safe against discarding signals from our target population (see next Section). So thousands of test signals from the target population are added to the data, each is searched for, exactly as done for the real search, and the corresponding $R^{a}$ is measured. The threshold dismisses at each stage less than $0.01\%$ of the test signals, resulting in an overall false dismissal of less than $0.04\%$.
@@ 6.2 Target population
The post-processing procedures that follow the Einstein@Home search have been designed to have a very low false dismissal for signals of the target population. The target population signals for a given supernova remnant obviously all come from the position of that source and have frequency and frequency derivative values in the target range, randomly displaced from the search grid points. The $\cos\iota$ and $\psi$ parameters are uniformly distributed $-1\leq\cos\iota\leq 1$, $-\pi/4\leq\psi\leq\pi/4$. The signal strength is set to the target strength with respect to the amplitude spectral density: the target population signals have amplitudes corresponding to $\mathcal{D}=90~{}[1/\sqrt{\textrm{Hz}}]$ for both Vela Jr. and Cas A (see Eq. 15 for the relation between the amplitude $h_{0}$ and $\mathcal{D}$). The target depth value is the best (i.e. highest sensitivity) that we could achieve with the available computational power, and in absolute terms it is the highest achieved on such a large parameter space.
@@ 6.3 Stage-0
The goal at this stage is to efficiently select the most compelling candidates from the $5.7\times 10^{10}$ results returned to the Einstein@Home server. Historically, there have been three main ways to achieve this.
@@
The third method is called clustering. In clustering, candidates associated with the same astrophysical or non-astrophysical causes are identified, and one of them is chosen (A. Singh et al., 2017; B. Steltner et al., 2022a). Clustering is generally efficient because it prevents the waste of resources on following up candidates originatign from the same root cause. While clustering has proven to be very effective in all-sky searches, in directed searches, the sensitivity improvement compared with simple thresholding is very modest – less than $1$ (J. Ming et al., 2024). Since determining a good operating point for the clustering procedure requires time and extensive Monte Carlos, we do not pursue it here. Instead, we propose an alternative scheme to select the Stage-0 results to follow-up: the pixeling procedure. This alternative is straightforward and efficient and does not loose detection efficiency in disturbed frequency bands.
@@ 6.3.1 The Pixeling Procedure
We divide the searched frequency range in $N_{\textrm{bands}}$ bands of equal size. For each frequency band, we consider $N_{p}$ frequency-spindown “pixels”, with each pixel being a unique small region of the searched parameter space. The frequency range covered by each pixel is 5 mHz. The $\dot{f}$ range covered by a pixel varies with frequency and with the age $\tau$ of the source, and it is equal to $f/10\tau$. So we consider 0.5 Hz wide frequency bands, and $N_{p}=100\times 10$ pixels to cover the $f\times\dot{f}$ parameter space.
We take the top $n_{px}=2$ Stage-0 results from waveforms with parameters in each pixel, with $n_{px}$ determined by the available computational budget for the follow-up. We call these candidates. The lowest detection statistic in the pixels from a 50 mHz region is shown in Figure 2 as an example.
In all we select $\approx 1.5$ million candidates from the Cas A search and $\approx 5.2$ million from Vela Jr. search. Both these numbers are $<2N_{p}\times N_{\textrm{bands}}$ (twice the total number of pixels) because not all pixels contribute 2 candidates – in fact some contribute no candidate at all. This is due to the fact that the original results are themselves top-lists returned by the Einstein@Home volunteer computers, and in the presence of very large disturbances most results might be concentrated in certain frequency-spindown regions rather than being $\approx$ uniformly distributed among the pixels. The average value of the $\hat{\beta}_{\text{S/GLtLr}}$ detection statistic of the considered candidates increases by 4 going from 20 to 400 Hz for Cas A and increases by 2 going from 400 to 1700 Hz for Vela Jr. This is due to the a trials factor effect: we are taking the most significant 2 results from a parameter space (the pixel) that increases in size as the frequency increases.
@@
Figure 2: An example of the lowest recorded $\hat{\beta}_{\text{S/GLtLr}}$ per pixel in a slightly disturbed (top-left) band. The results shown come from the Vela Jr. search. Each cell in this figure is a “pixel”.
@@ 7.1 Intrinsic Gravitational Wave Amplitude Upper Limits
The detection efficiency is the number of recovered signals relative to the total number of signals considered. Each test signal is added to the data and then the with the test signals is treated exactly like the data from original Einstein@Home search, including data preparation, cleaning and candidate-selection with pixeling. A signal is considered recovered if a result from the search with the test signal, and ascribable to the signal, survives until the last stage. In practice, since the confidence in recovering signals at our upper limit amplitudes during the follow-up stages is much higher (99.96%) than the upper limit confidence level ($90\%$) we neglect the false dismissal incurred by the follow-up and significantly simplify the upper limit pipeline by requiring that the test signal only survive Stage-0.
As described in Section 4, in frequency bands affected by large spectral-line disturbances, we substitute the real data with Gaussian noise. For consistency we also do this after the test signal has been added to noise, hence, depending on the data removed ands on the frequency of the test signal, it might not be possible to reach the 90% confidence level for any value of $h_{0}$. For frequency bands where this happens we do not provide upper limits. We add these bands to the list of excluded bands given in J. Ming et al. (2025). Such bands include 6 half-Hz bands between 20 Hz and 400 Hz and 109 half-Hz bands between 400 Hz and 1700 Hz search, mostly due to the thermal excitation of vibration modes of the suspension silica fibers at 500 Hz and their harmonics around 1000 Hz and 1500 Hz (B. P. Abbott et al., 2016).
@@ 7.2 Upper Limits Recast
$\alpha=0.028\left(\frac{h_{0}}{10^{-24}}\right)\left(\frac{D}{1\ \text{kpc}}\right)\left(\frac{100\ \text{Hz}}{f}\right)^{3}.$ (14)
@@ 7.3 Sensitivity Depth
The average sensitivity depth across the frequency ranges is approximately $92[{1/\sqrt{\text{Hz}}}]$ for Cas A and $89[{1/\sqrt{\text{Hz}}}]$ for Vela Jr., respectively. Although the average mismatch of the original Einstein@Home search for the Vela Jr. is slightly smaller than for Cas A, the average sensitivity depth of search for Vela Jr. is slightly larger. The reason for this is that the follow-ups involve proportionally about 3.5 times more candidates for Cas A than for Vela Jr.
@@ 8 Conclusions
We present results for searches for continuous gravitational waves from the neutron star in the Cas A and Vela Jr. supernova remnants. The considered waveforms cover frequencies in the range 20 - 400 Hz for Cas A and 400 - 1700Hz Vela Jr., and very broad frequency evolution scenarios including power-law frequency evolutions with braking indexes up to 7. In particular, the second-order spin-down range of our searches is $\ddot{f}\in[0,7{f/\tau^{2}}]$, which is broader than the range covered by R. Abbott et al. (2022d) and J. Wang & K. Riles (2024), who assume $[2|\dot{f}|^{2}/f,7|\dot{f}|^{2}/f]$. Furthermore, our search range is independent of the template $\dot{f}$ value, accommodating signals that deviate from a strict power law.
The large parameter space, comprising in all $1.3\times 10^{18}$ waveforms, could be investigated thanks to the computing power donated by the volunteers of the Einstein@Home project. Promising candidates across this space were identified using a new “pixeling” method and followed-up in a cascade of further searches. Fewer and fewer candidates survive each stage and at the end of the fourth stage no signal candidate is left, indicating no detectable signal.